OpenStudy (anonymous):

Two square tables pushed together makes a rectangle that has a surface area of 338 inches squared. How long is each side of a square table ?

OpenStudy (gudden):

alright so are we supposed to find the length of the side of the square ??

OpenStudy (anonymous):

Yeah sorry didn't put that

OpenStudy (gudden):

 and area of rectangle = length * breadth here length = 2x and breadth = x on multiplying length with breadth we get 2x^2 = area of rectangle and area of rectnagle has been given as 338 so lets equate these 2x^2 = 338 on dividing both sides of equation with 2 get x^2 = 169 and on putting a square root we get x = 13 so sides of square are 13 inches